The Organizing Committee of the 15-th ILAS Conference thanks

Albrecht Boettcher

TU Chemnitz, Chemnitz, Germany

for delivering the Invited Plenary Talk “Toeplitz matrices with Fisher-Hartwig symbols".

Heike Fassbender

TU Braunschweig, Braunschweig, Germany

for delivering the Invited Plenary Talk

“Structured Methods for Eigenproblems with Hamiltonian Spectral Symmetry" (with Peter Benner and Martin Stoll).

The Organizing Committee of the 15-th ILAS Conference thanks

Luca Gemignani

Department of Mathematics University of Pisa, Pisa, Italy

for delivering the Invited Plenary Talk

“Eigenvalue Problems for Rank-structured Matrices".

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Leslie Hogben

Iowa State University, Ames, Iowa, USA for delivering the Invited Plenary Talk

“Minimum Rank Problems: Recent Developments".

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

Er-Xiong Jiang

Shanghai University, Shanghai, China for delivering the Invited Plenary Talk

“Some inverse eigenvalue problems for Jacobi matrices".

Luis Verde-Star Chair of the Organizing Committee

Daniel Kressner

ETH Zurich, Zurich, Switzerland

for delivering the Invited Plenary Talk “Matrix product eigenvalue problems".

The Organizing Committee of the 15-th ILAS Conference thanks

Froilán M. Dopico

Universidad Carlos III, Madrid, Spain for delivering the Invited Plenary Talk

“Implicit Jacobi algorithms for the symmetric eigenproblem".

Luis Verde-Star Chair of the Organizing Committee The Organizing Committee of the 15-th ILAS Conference thanks

James Nagy

Emory University, Atlanta, USA for delivering the Invited Plenary Talk “Kronecker Products in Imaging Sciences".

Luis Verde-Star Chair of the Organizing Committee

Juan Manuel Peña

University of Zaragoza, Zaragoza, Spain

for delivering the Invited Plenary Talk

“From Total Positivity to Positivity: related classes of matrices".

Peter Rosenthal

University of Toronto, Canada

for delivering the Invited Plenary Talk “Invariant subspaces of semigroups of matrices".

Naomi Shaked-Monderer

Emek Yezreel College, Emek Yezreel, Israel

for delivering the Invited Plenary Talk “Completely Positive Matrices and the CP-rank".

The Organizing Committee of the 15-th ILAS Conference thanks

Ilya Spitkovsky

College of William and Mary, Blacksburg, VA, USA for delivering the Invited Plenary Talk

“On the current state of the factorization problem for almost periodic matrix functions".

Luis Verde-Star Chair of the Organizing Committee

Paul Van Dooren

Université Catholique de Louvain, Louvain-la-Neuve, Belgium

for delivering the Invited Plenary Talk

“Some graph optimization problems in data mining".